# The Anderson transition – a view from Krylov space

### v1

@@ Abstract
The Krylov subspace expansion is a workhorse method for sparse numerical methods that has been increasingly explored as source of physical insight into many-body dynamics in recent years. We revisit the venerable Anderson model of localization in dimensions $d=1,2,3,4$ to construct local integrals of motion (LIOM) in Krylov space. These appear as zero eigenvalue edge states of an effective hopping problem in Krylov superoperator subspace, and can be analytically constructed, given the Lanczos coefficients. We exploit this idea, focusing on $d=3$, to study the manifestation of the disorder driven Anderson transition in the anatomy of LIOMs. We find that the increasing complexity of the Krylov operators results in a suppression of the fluctuations of the Lanczos coefficients. As such, one can study the phenomenology of the integrals of motion in the disorder averaged Krylov chain. We find edge states localized on vanishing fraction of Krylov space (of dimension $D_{K}=V^{2}$ for cubes of volume $V$), both in localized and extended phases. Importantly, in the localized phase, disorder induces powerlaw decaying dimerization in the (Krylov) hopping problem, producing stretched exponential decay of the LIOMs (in Krylov space) with a stretching exponent $1/2d$. Metallic LIOMs are completely delocalized albeit across only $\propto\sqrt{D_{K}}$ states). Critical LIOMs exhibit powerlaw decay with an exponent matching the expected fractal exponent.
@@ I Introduction
Understanding the effects of quenched disorder on electronic wave functions has been instrumental in understanding the electronic properties of solids. In the case of the non-interacting Anderson model, spatial disorder is known to localize the electronic states if it is larger than a critical disorder strength, whereas below this strength the electron states are extended throughout the system [1, 2, 3, 4]. Applying these concepts to interacting systems has been a major endeavor in the past decades [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 8, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 28], but the interacting problem, dubbed many-body localization (MBL), has proved significantly more challenging both conceptually and numerically [35, 36, 37, 38].
@@
We understand these localized ergodicity-breaking systems to be effectively integrable, possessing an extensive set of local integrals of motion (LIOMs). One may motivate their existence intuitively, invoking adiabatic evolution of the Hamiltonian from completely decoupled degrees of freedom [39], e.g. qubits, in which case LIOMs are Pauli-like operators, or they can be obtained by simply time-averaging Hamiltonian evolution of initially localized qubits [40, 41]. LIOMs must be exponentially localized with a small enough localization length to be stable against so called “quantum avalanches” induced by rare Griffith’s regions [42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 17, 52, 53, 54, 55]. Given the non-uniqueness and the stringent conditions on their decay, it is important to continue looking for new methods to understand the emergence of local integrals of motion. The Krylov subspace expansion, commonly known as the recursion method [56], is one such method which, although not a new technique, has been increasingly explored as a probe of quantum chaos (for a recent review, see Ref. [57]). The LIOMs explored in this work are the same operators as the time-averaged LIOMs of Ref. [40], but, importantly, expressed in the dynamically generated Krylov basis, which is arguably more natural and demonstrably more compact than any apriori fixed basis.
@@
The edge mode nature of LIOMs studied below has been identified in recent work on nearly integrable spin chains [58, 59, 60]. By way of surveying other recent efforts to exploit Krylov methods, we note studies of quantum chaos and integrability in Krylov space [61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 59, 60, 78, 79, 80, 81, 58, 82, 83, 84, 85, 86], and proposals for detecting ergodicity breaking transitions. For one, chaos has been associated with a maximum growth rate of the Lanczos coefficients [61], though this is not always the case [68]. The interplay between integrability, integrability breaking, and dimerization has been pointed out in Refs. [64, 71, 57]. Finally, chaos is associated with delocalization in Krylov space [87], which can be quantified by the average position in Krylov space known as the “Krylov complexity”, the long time average of which is expected to saturate a maximum bound for chaotic systems [61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 73].
@@
In the balance of the paper we introduce the operator Krylov space, generated from a chosen probe operator, and show how the Lanczos coefficients contain all of the information of that operator’s dynamics (in Section II). Section II.1 explains how to construct integrals of motion in Krylov space. Section III introduces the Anderson model and presents some summary statistics of its Lanczos coefficients for a particular probe, clearly identifying three different asymptotic decays of zero-modes in the localized phase, critical point, and extended phase. We close (Sec. IV) with a summary of results and open problems.
@@ II Operator Krylov space
We will now briefly describe the Krylov subspace expansion which is explained in greater detail in Refs. [56, 57]. In the Heisenberg picture operators $O$ evolve in time under the action of a Hamiltonian $H$ as $\frac{dO}{dt}=i[H,O]$, which upon defining the Liouvillian $\mathcal{L}(\cdot)=[H,\cdot]$, allows for the formal solution $O(t)=e^{i\mathcal{L}t}(O)$. One can do a Taylor expansion of the latter which illuminates the basic idea of the recursion method: to recursively construct a basis of orthogonal polynomials. To define these requires defining an inner product on the operator space, and this choice is not unique. Since we’re concerned with eigenstate transition at arbitrary energy, we choose the (normalized) Hilbert-Schmidt norm $O||^{2}=D^{-1}{\rm Tr}[O^{\dagger}O]$, where $D$ is the dimension of Hilbert space.
@@
Therefore the Lanczos coefficients $\{b_{n}\}$, which have units of energy, contain all of the information needed to compute the dynamics of the desired operator $O_{1}$. The dimension of Krylov space $D_{K}$ is related to the number of distinct frequencies and thus depends on the model being considered, however it is upper bounded at $D_{K}\leq D^{2}-D+1$ [63]. Systems with level-repulsion are expected to saturate this bound, and although some special integrable systems have much smaller $D_{K}$ [68]: these systems have a small dynamical Lie algebra. In absence of those dynamical symmetries the Krylov space is typically maximal and the long time limit of the average position in Krylov space, otherwise known as the Krylov complexity, is often used as a measure of integrability [57]).
In addition, as the Lanczos coefficients $b_{n}$} are calculated through repeated action of $n$ powers of $\mathcal{L}$, they are equivalent to knowing the moments of the Liouvillian $\mu_{2n}$, which is explained in detail in Ref. [56]. The autocorrelation function is also simply defined in this basis as:
$C(t)=\frac{1}{D}{\rm Tr}\left[O_{1}e^{i\mathcal{L}t}(O_{1})\right],$ (3)
Aside from numerical considerations, this construction has conceptual advantage. We can consider $L$ as a Hamiltonian of a different system, which due to its tri-diagonal nature corresponds to a single particle hopping on a semi-infinite chain, with $b_{n}$} as the hopping amplitudes. In particular, if we express the operator as
@@
In this work we focus on calculating the integrals of motion, or conserved quantities, in Krylov space. Conserved quantities $Q$ commute with the Hamiltonian, i.e. $\mathcal{L}(Q)=0$ which in Krylov space corresponds to solutions to equation (5) with $\partial_{t}\phi_{n}=0$. These local integrals of motion would depend on the initial operator $O_{1}$. In particular, for a local seed operator $O_{1}$ and for local Hamiltonians the Krylov expansion is local, i.e. the range of the operator expands by the range of the interactions in the Hamiltonian in each iteration. This offers the interesting possibility to obtain integrals of motion $Q$ that have particularly high overlap with some desired operator $O_{1}$.
@@ II.1 Krylov space integrals of motion
The tri-diagonal structure of $L$ gives a simple solution to the zero-mode as a function of $n$ along the semi-infinite chain, in terms of the initial weight on the first site $Q_{1}$. The support on all the even $n$ vanishes, as they describe currents, while for odd $n$ one arrives at following form:
@@
In the ideal case, we’d get an exponentially localized zero-mode on the boundary of our Krylov chain, i.e. $\kappa(n)\sim n$. The latter corresponds to a perfectly dimerized chain in which all the odd couplings are greater than the even couplings. This situation essentially describes the Su–Schrieffer–Heeger (SHH) chain [89] in the topologically non-trivial insulating phase, and our integral of motion corresponds to the well known edge state. While this scenario is realized for the Majorana zero-mode in the transverse field Ising (TFI) chain [58, 59], it’s extremely fine tuned and unstable to small perturbations [60]. There are some generic properties that highlight the difficulty of finding these types of integrals of motion.
First of all, generic seed operators $O_{1}$ are expected to have matrix elements between all eigenstates of $H$ and we would therefore not expect them to have a spectral gap, e.g. in the eigenstates thermalization (ETH [88]) regime we expect low frequency matrix elements (below the Thouless scale) to behave like random matrix theory. The SSH scenario outlined above is thus highly non-generic as it corresponds to a problem in which our initial operator $O_{1}$ has a gapped spectrum. Generically, we’d need a scenario where there is localization on the Krylov chain but the spectrum is gapless. Random, or some form of pseudo-random, Lanczos coefficients could give us that [90], but as we’ll discuss later this turns out to have other problems. The obvious issue is that, while this results in localization, it would be hard to guarantee that the zero-mode is localized close to the boundary.
For local models the Lanczos coefficients are bounded to grow at most linearly $(b_{n}\sim n)$, with a logarithmic correction in 1D $(b_{n}\sim n/\log(n))$ [61, 74, 57]. Typically this growth is saturated and it has the tendency to cause exponentially fast transport on the Krylov chain, further hampering the construction of integrals of motion.
In order to disentangle the delocalizing effects of many-body operator growth from the localizing effects of disorder, we focus our study on the non-interacting Anderson model. The latter has the additional benefit that we can numerically access much larger system sizes.
@@ III Single particle localization in Krylov space
Here $c^{\dagger}_{i},c_{j}$ are fermionic creation and annihilation operators on sites $i$ and $j$ respectively, the first sum runs over nearest neighbors $\langle i,j\rangle$, and the second sum runs over all sites $L^{d}$ where $d$ is the spatial dimension. The particle has a hopping amplitude $t$, which we set to $t=1$, and a random on-site potentials which are constructed from i.i.d. uniformly distributed random numbers $h_{i}\in[-1,1]$, we refer to $W$ as the disorder strength. In one and two spatial dimensions this model is localized for any $W$, while in three dimensions and higher the model is initially delocalized at low disorder, but undergoes an infinite temperature metal-insulator phase transition at a critical disorder, which is estimated at $W_{c}\approx 16.5$ in three dimensions [91, 92, 93].
Figure 1: Normalized variance of the Lanczos coefficients for the three-dimensional Anderson model at $W=4,16,40$. The inset shows an anti-diagonal cut of the correlation matrix $\rho_{ij}={\rm cov}[b_{i},b_{j}]/\sigma(b_{i})\sigma(b_{j})$ which we denote $\rho_{i,N+1-i}$ and which is sharply peaked only at the diagonal index ($i=0$ here). Different color shades indicate different system sizes; from lightest to darkest are $L=14,18,22$. All results have been obtained from $N_{s}=1000$ samples.
Figure 2: Disorder averaged Lanczos coefficients of the three-dimensional Anderson model over a range of different disorders $W$, both above and below the critical disorder. Results are obtained for a system size of $L=22$ with $N_{s}=1000$ samples each.
Because we are ultimately interested in constructing the conserved quantities, which in the Anderson model are dressed local density operators, we begin our Krylov expansion with a local traceless single site density operator: $O_{1}=c^{\dagger}_{x}c_{x}-1/2$. Here the site $x$ is the site around which we’d like to construct our LIOM. We can then write the Krylov expansion of the many-particle problem completely in the single-particle operators, since the entire problem is quadratic in fermionic operators. The single particle Hamiltonian $h$ (an $L^{d}\times L^{d}$ matrix) is defined as $H=\sum_{ij}h_{ij}c^{\dagger}_{i}c_{j}$. The single-particle Green function $G$ allows us to express the Heisenberg evolved operators as:
$c_{x}(t)=\sum_{y}c^{\dagger}_{y}G(y,t|x,0)=\sum_{y}c^{\dagger}_{y}\left<y|e^{-iht)}|x\right$>. (9)
By Wick’s theorem we arrive at following expression for the autocorrelation $C(t)$:
@@
which is just $1/4$ of the return probability to site $x$. Note that it was important to start with a traceless operator. In conclusion we can simply perform the Krylov expansion on the single-particle problem by taking a Hilbert-Schmidt inner product on the single particle space and identifying the initial many-body operator $c^{\dagger}_{x}c_{x}-1/2$ with the single particle projector $P_{x}=1/2\left|x\right>\left<x\right|$.
@@
The mean Lanczos coefficients are shown in Fig. 2, where a clear initial dimerization is observed, which then decreases by some amount. The average $b_{n}$ across $n$ grows slightly at first but quickly saturates as expected for a free system. In the extended regime (see e.g. $(W=10)$ in Fig. 2) this dimerization is very small, but as we will show next, it still affects the form of the zero-mode.
Figure 3: Parameters of the dimerization model, expression 11, fitted by least squares as a function of disorder (lighter dots show the confidence interval at $10\%$ significance). The critical point of the model, corresponding to $\gamma=0$, occurs at $W_{c}\approx 16$, consistent with the accepted value $W_{c}\simeq 16.5$. The black dotted line marks the transition point $\gamma=0$ and the grey dotted line marks the localized prediction $\gamma=1/2d$. In the inset, the grey dotted line marks the predicted value in the extended regime: $\epsilon=L^{-d}$
Since fluctuations are small, we focus our attention on the disorder averaged coefficients, allowing us to define a simple model for the dimerization of the coefficients:
@@
We can extract the model parameters in (11) by fitting it to the coefficients shown in Fig. 2. The results for the parameters, using a least squares fitting procedure, are shown in Fig. 3. We find $\gamma\leq 0$ at weak disorder and crosses zero around $W\approx 16$, close to the accepted $W_{c}\approx 16.5$.

### v2

@@ Abstract
The Krylov subspace expansion is a workhorse method for sparse numerics that has been increasingly explored as source of physical insight into many-body dynamics in recent years. In this work we revisit the venerable Anderson model of localization in dimensions $d=1,2,3,4$ to construct local integrals of motion (LIOM) in Krylov space. These appear as zero eigenvalue edge states of an effective hopping problem in the Krylov superoperator subspace and can be analytically constructed given the Lanczos coefficients. We exploit this idea, focusing on $d=3$, to study the manifestation of the disorder driven Anderson transition in the anatomy of LIOMs. We find that the increasing complexity of the Krylov operators results in a suppression of the fluctuations of the Lanczos coefficients. As such, one can study the phenomenology of the integrals of motion in the disorder averaged Krylov chain. We find edge states localized on vanishing fraction of Krylov space (of dimension $D_{K}=V^{2}$ for cubes of volume $V$), both in localized and extended phases. Importantly, in the localized phase, disorder induces powerlaw decaying dimerization in the (Krylov) hopping problem, producing stretched exponential decay of the LIOMs in Krylov space with a stretching exponent $1/2d$. Metallic LIOMs are completely delocalized albeit across only $\propto\sqrt{D_{K}}$ states. Critical LIOMs exhibit powerlaw decay with an exponent matching the expected value of $0.29$.
@@ I Introduction
Modeling quenched disorder has been instrumental in understanding the electronic properties of solids. In the case of the non-interacting Anderson model, spatial disorder is known to localize the electronic states if it is larger than a critical disorder strength, whereas below this strength the electron states are extended throughout the system [Anderson, Mott1968-3d, Kramer1993-localization_review, Evers_2008_AndersonTransitions]. Applying these concepts to interacting systems has been a major endeavor in the past decades [BASKO20061126, gornyi2005MBL, Oganesyan2007, annurev-conmatphys-031214-014726, Huse2014Phenomenoloy, Imbrie2016, ROS2015420, Serbyn2013, Thomson2018, AbaninReview, annurev-conmatphys-031214-014726, ALET2018498, Schreiber2015, Rubio-Abadal2019, Smith2016, Kohlert2019, Lukin2019, Rispoli2019, Pal2010MBLTransision, Lev2015, Luitz2015, Wahl2017, Mierzejewski2016, Mace2019, Chanda2020, Laflorencie2020, Gray2018, Doggen2018, POLFED2020, Serbyn2015MBLTransition, Sierant2020ThoulessTime, Chanda2020], but the interacting problem, dubbed many-body localization (MBL), has proved significantly more challenging both conceptually and numerically [suntajs2019quantum, suntajs2020transition, Dynamicalobstruction, sierant2025mbl].
@@
We may interpret these localized ergodicity-breaking systems to be effectively integrable, meaning they must possess an extensive set of local integrals of motion (LIOMs). Their existence may be motivated intuitively, invoking adiabatic evolution of the Hamiltonian from completely decoupled degrees of freedom [Imbrie2017_LIOMs], e.g. qubits, in which case LIOMs are Pauli-like operators (i.e. having discrete spectra $\pm 1$), or they can be obtained by simply time-averaging the Hamiltonian evolution of initially localized qubits [chandran2015constructing, geraedts2017emergent]. Though not uniquely defined, LIOMs should be exponentially localized with a small enough localization length to be stable against so called “quantum avalanches” induced by rare Griffith’s regions [Chandran2016_MBLbeyondEigenstates, deRoeck17, Luitz2017_HowaSmallQuantumBathCanThermalize, Ponte2017_HowOneSpin, MorningstarAvalanches, Bath-induceddelocalization, Embedded_Thermal_inclusion, Selsthermalizationofdiluteimpurities, Crowley2020_AvalancheInducedLocalizedandThermalRegions, Potirniche2019_ExplorationofStabilityofMBL, Rubio-Abadal2019, Crowley2022_MeanFieldTheoryofFailedThermalizingAvalanches, Suntajs2022_ErgodicityBreakinginZeroDimensions, Long2023_Prethermal_Phenom, Leonard2023avalanchesexperiment]. Given their non-unique construction and the stringent conditions on their decay, it is important to continue looking for new methods to understand the emergence of local integrals of motion. The Krylov subspace expansion, commonly known as the recursion method [recursionmethod_vis], is one such method which, although not a new technique, has been increasingly explored as a probe of quantum chaos (for a recent review, see Ref. [nandy2024quantumdynamicskrylovspace]). The LIOMs we will explore in this work are the same operators as the time-averaged LIOMs of Ref. [chandran2015constructing], but, importantly, expressed in the dynamically generated Krylov basis, which is arguably more natural and demonstrably more compact than any apriori fixed basis.
@@
The edge mode nature of LIOMs studied below has been identified in recent work on nearly integrable spin chains [Yates2020_LifetimeofAlmostStrong, Yates2021StrongandAlmoststrongModes, Yeh2023SlowlyDecayingZeroMode]. By way of surveying other recent efforts to exploit Krylov methods, we note there have been many studies of quantum chaos and integrability in Krylov space [Parker2019_universaloperatorgrowth, Barbón2019OnTheEvolutionofOperatorComplexity, Rabinovici2021_OperatorComplexityaJourney, Rabinovici2022Krylovlocalizationandsuppression, Hornedal2022_UltimateSpeedLimitstoGrowthofOpComplexity, Caputa2022_GeometryofKrylovComplexity, Balasubramanian2022_QuantumChaosandtheComplexityofSpreadofStates, Bhattacharjee2022_KComplexity_SaddleDominatedScrambling, Erdmenger2023_UniversalChaoticDynamicsfromKrylov, Rabinovici2022_KrylovComplexityfromIntegrabilitytoChaos, Hashimoto2023_KrylovComplexityandChaos, Suchsland2025_KrylovComplexityandTrotterTransitions, Camargo2024_SpectralandKrylovComplexityBilliards, Cao_2021StatisticalmethodforOG, menzler2024krylovlocalizationprobeergodicity, baggioli2024_krylovcomplexityorderparameter, Yates2020DynamicsofAlmostStrongEdgeModes, Yates2021StrongandAlmoststrongModes, Yeh2023SlowlyDecayingZeroMode, Yeh2024UniversalModelofFloquet, loizeau2025openingkrylovspaceaccess, Trigueros2022_KrylovComplexityofMBL, Balasubramanian2025_ChaosIntegrabilityandLateTimes, Yates2020_LifetimeofAlmostStrong, Scialchi2025_ExploringQuantumErgodicity_KrylovApproach, Uskov2024_QuantumDynamicsinOneandTwoDims, Teretenkov_2025_PseudomodeExpansion, Loizeau_2025_PauliStrings, Yeh2025_MomentMethodandContinuedFractionExpansion] which come with various proposals for detecting ergodicity breaking transitions. For one, chaos has been associated with a maximum growth rate of the Lanczos coefficients [Parker2019_universaloperatorgrowth], though this is not always the case [Bhattacharjee2022_KComplexity_SaddleDominatedScrambling]. The interplay between integrability, integrability breaking, and dimerization has been pointed out in Refs. [Rabinovici2022Krylovlocalizationandsuppression, Hashimoto2023_KrylovComplexityandChaos, nandy2024quantumdynamicskrylovspace, menzler2024krylovlocalizationprobeergodicity]. Finally, chaos is associated with delocalization in Krylov space [Dymarsky2020_QuantumChaosAsDelocalization], which can be quantified by the average position in Krylov space known as the “Krylov complexity”, the long time average of which is expected to saturate a maximum bound for chaotic systems [Parker2019_universaloperatorgrowth, Barbón2019OnTheEvolutionofOperatorComplexity, Rabinovici2021_OperatorComplexityaJourney, Rabinovici2022Krylovlocalizationandsuppression, Hornedal2022_UltimateSpeedLimitstoGrowthofOpComplexity, Caputa2022_GeometryofKrylovComplexity, Balasubramanian2022_QuantumChaosandtheComplexityofSpreadofStates, Bhattacharjee2022_KComplexity_SaddleDominatedScrambling, Erdmenger2023_UniversalChaoticDynamicsfromKrylov, Rabinovici2022_KrylovComplexityfromIntegrabilitytoChaos, Hashimoto2023_KrylovComplexityandChaos, Camargo2024_SpectralandKrylovComplexityBilliards].
@@
In the balance of the paper we introduce the operator Krylov space, generated from a chosen probe operator, and show how the associated Lanczos coefficients can be used to both reproduce that operator’s dynamics and construct LIOMs (Section II). The Anderson model is introduced in Section III and we present some summary statistics of its Lanczos coefficients for a particular probe, clearly identifying three different asymptotic decays of zero-modes in the localized phase, critical point, and extended phase. We close (Sec. IV) with a summary of results and open problems.
@@ II Operator Krylov space
We will now briefly describe the Krylov subspace expansion which is explained in greater detail in Refs. [recursionmethod_vis, nandy2024quantumdynamicskrylovspace]. In the Heisenberg picture operators $O$ evolve in time under the action of a Hamiltonian $H$ as $\frac{dO}{dt}=i[H,O]$, which upon defining the Liouvillian $\mathcal{L}(\cdot)=[H,\cdot]$, allows for the formal solution $O(t)=e^{i\mathcal{L}t}(O)$. One can do a Taylor expansion of the latter which illuminates the basic idea of the recursion method: to recursively construct a basis of orthogonal operator polynomials. To define these requires defining an inner product on the operator space, and this choice is not unique. Since we’re concerned with eigenstate transitions at arbitrary energy, we choose the (normalized) Hilbert-Schmidt norm $O||^{2}=D^{-1}{\rm Tr}[O^{\dagger}O]$, where $D$ is the dimension of the many-body Hilbert space.
@@
Therefore the Lanczos coefficients $\{b_{n}\}$, which have units of energy, contain all of the information needed to compute the dynamics of the desired operator $O_{1}$. The dimension of Krylov space $D_{K}$ is related to the number of distinct frequencies and thus depends on the model being considered, however it is upper bounded at $D_{K}\leq D^{2}-D+1$ [Rabinovici2021_OperatorComplexityaJourney]. Systems with level-repulsion are expected to saturate this bound, but some special integrable systems have much smaller $D_{K}$ [Bhattacharjee2022_KComplexity_SaddleDominatedScrambling] due to their smaller dynamical Lie algebra. In absence of those dynamical symmetries the Krylov space is typically maximal, and as previously noted, the long time limit of the average position in Krylov space, otherwise known as the Krylov complexity, is often used as a measure of integrability [nandy2024quantumdynamicskrylovspace]).
In addition, as the Lanczos coefficients $b_{n}$} are calculated through repeated action of $n$ powers of $\mathcal{L}$, they are equivalent to knowing the moments of the Liouvillian $\mu_{2n}$ (explained in detail in Ref. [recursionmethod_vis]). The autocorrelation function is also simply defined in this basis as:
$C(t)=\frac{1}{D}{\rm Tr}\left[O_{1}e^{i\mathcal{L}t}(O_{1})\right]$ (3)
Aside from numerical considerations, this construction has conceptual advantage. We can consider $\mathcal{L}$ as a Hamiltonian in its own right, which due to its tri-diagonal nature corresponds to a single particle hopping on a semi-infinite chain, with $b_{n}$} as the hopping amplitudes. In particular, if we express the operator as
@@
In this work we focus on calculating the integrals of motion, or conserved quantities, in Krylov space. Conserved quantities $Q$ commute with the Hamiltonian, i.e. $\mathcal{L}(Q)=0$, which in Krylov space corresponds to the zero-modes of $L$, which are solutions to expression (5) with $\partial_{t}\phi_{n}=0$. These local integrals of motion depend on the initial operator $O_{1}$. In particular, for a local seed operator $O_{1}$ and for local Hamiltonians the Krylov expansion is local, i.e. the range of the operator expands by the range of the interactions in the Hamiltonian in each iteration. This offers the interesting possibility to obtain integrals of motion $Q$ that have particularly high overlap with some desired operator $O_{1}$.
@@ II.1 Krylov space integrals of motion
The tri-diagonal structure of $\mathcal{L}$ in the Krylov basis gives a simple solution to the zero-mode as a function of $n$ along the semi-infinite chain in terms of the initial weight on the first site $Q_{1}$. The support on all the even $n$ vanishes, as they describe currents, while for odd $n$ one arrives at following form:
@@
In the ideal case, we’d get an exponentially localized zero-mode on the boundary of our Krylov chain, i.e. $\kappa(n)\sim n$. The latter corresponds to a perfectly dimerized chain in which all the odd couplings are greater than the even couplings. This situation essentially describes the Su–Schrieffer–Heeger (SHH) chain [SSH1979] in the topologically non-trivial insulating phase, and our integral of motion corresponds to the well known edge state. While this scenario is realized for the Majorana zero-mode in the transverse field Ising (TFI) chain [Yates2020_LifetimeofAlmostStrong, Yates2021StrongandAlmoststrongModes], it’s extremely fine tuned and unstable to small perturbations [Yeh2023SlowlyDecayingZeroMode].
There are some generic properties that highlight the difficulty of finding these types of integrals of motion. First of all, generic seed operators $O_{1}$ are expected to have matrix elements between all eigenstates of $H$ and we would therefore not expect them to have a spectral gap, e.g. in the eigenstates thermalization regime (ETH [Srednicki1994ETH]) we expect low frequency matrix elements (below the Thouless scale) to behave like random matrix theory. The SSH scenario outlined above is thus highly non-generic as it corresponds to a problem in which our initial operator $O_{1}$ has a gapped spectrum. Generically, we’d need a scenario where there is localization on the Krylov chain but the spectrum is gapless. Random, or some form of pseudo-random, Lanczos coefficients could give us that [Fleishman1977_FluctuationsandLocalizationin1D], but as we’ll discuss later this turns out to have other problems. The primary issue is that, while this results in localization, it would be hard to guarantee that the zero-mode is localized close to the boundary. In addition, for local models the Lanczos coefficients are bounded to grow at most linearly $(b_{n}\sim n)$, with a logarithmic correction in 1D $(b_{n}\sim n/\log(n))$ [Parker2019_universaloperatorgrowth, Cao_2021StatisticalmethodforOG, nandy2024quantumdynamicskrylovspace]. Typically this growth is saturated and it has the tendency to cause exponentially fast transport on the Krylov chain, further hampering the construction of integrals of motion.
In order to disentangle the delocalizing effects of many-body operator growth from the localizing effects of disorder, we focus our study on the non-interacting Anderson model. This model has the additional benefit that we can numerically access much larger system sizes.
@@ III Single particle localization in Krylov space
Here $c^{\dagger}_{i},c_{j}$ are fermionic creation and annihilation operators on sites $i$ and $j$ respectively, the first sum runs over nearest neighbors $\langle i,j\rangle$ with periodic boundary conditions, and the second sum runs over all sites $L^{d}$ where $d$ is the spatial dimension. The particle has a hopping amplitude $t$, which we set to $t=1$, and random on-site potentials which are constructed from i.i.d. uniformly distributed random numbers $h_{i}\in[-1,1]$. We refer to $W$ as the disorder strength. In one and two spatial dimensions this model is localized for any $W$, while in three dimensions and higher the model is initially delocalized at low disorder, but undergoes an infinite temperature metal-insulator phase transition at a critical disorder, which is estimated at $W_{c}\approx 16.5$ in three dimensions [Corrections_to_Scaling_Anderson1999, Statistics_of_Spectra_Anderson1993, SUNTAJS_2021_Spectral_properties_3D].
Figure 1: Normalized variance of the Lanczos coefficients for the three-dimensional Anderson model at $W=4,16,40$. The inset shows an anti-diagonal cut of the correlation matrix $\rho_{ij}={\rm cov}[b_{i},b_{j}]/\sigma(b_{i})\sigma(b_{j})$ which we denote $\rho_{i,N+1-i}$ and which is sharply peaked only at the diagonal index ($i=0$). Different color shades indicate different system sizes: from lightest to darkest are $L=14,18,22$. All results have been obtained from $N_{s}=1000$ samples.
Figure 2: Disorder averaged Lanczos coefficients of the three-dimensional Anderson model over a range of different disorders $W$, both above and below the critical disorder (in order from top to bottom are $W=40,30,22,20,18,16,14,12,10,4$). All results are obtained for a system size of $L=22$ with $N_{s}=1000$ samples.
Because we are ultimately interested in constructing the conserved quantities, which in the Anderson model are dressed local density operators, we begin our Krylov expansion with a local traceless single site density operator: $O_{1}=c^{\dagger}_{x}c_{x}-1/2$. Here the site $x$ is the site around which we’d like to construct our LIOM. We can then write the Krylov expansion of the many-particle problem completely in the single-particle operators, since the entire problem is quadratic in fermionic operators. The single particle Hamiltonian $h$ (an $L^{d}\times L^{d}$ matrix) is defined as $H=\sum_{ij}h_{ij}c^{\dagger}_{i}c_{j}$. The single-particle Green function $G$ allows us to express the Heisenberg evolved operators as $c_{x}(t)=\sum_{y}c^{\dagger}_{y}G(y,t|x,0)=\sum_{y}c^{\dagger}_{y}\left<y|e^{-iht}|x\right$>. By Wick’s theorem we arrive at following expression for the autocorrelation $C(t)$:
@@
which is just $1/4$ of the return probability to site $x$. Note that it was important to start with a traceless operator. In conclusion, we can simply perform the Krylov expansion on the single-particle problem by taking a Hilbert-Schmidt inner product on the single particle space and identifying the initial many-body operator $c^{\dagger}_{x}c_{x}-1/2$ with the single particle projector $P_{x}=1/2\left|x\right>\left<x\right|$.
@@
The mean Lanczos coefficients are shown in Fig. 2, where a clear initial dimerization is observed, which then decreases by some amount. The average $b_{n}$ across $n$ grows slightly at first but quickly saturates as expected for a free system. In the extended regime (see e.g. $W=10$ in Fig. 2) this dimerization is very small, but as we will show next, it still affects the form of the zero-mode.
Figure 3: Parameters of the dimerization model, expression (10), fitted by a least squares procedure as a function of disorder (lighter dots show the confidence interval at $10\%$ significance). The critical point of the model, corresponding to $\gamma=0$, occurs at $W_{c}\approx 16$ consistent with the accepted value $W_{c}\simeq 16.5$. The black dotted line marks the transition point $\gamma=0$ and the grey dotted line marks the localized prediction $\gamma=1/2d$. In the inset, the grey dotted line marks the predicted value in the extended regime: $\epsilon=L^{-d}$
Now that we have established the fluctuations are small, we focus our attention on the disorder averaged coefficients, allowing us to define a simple model for the dimerization of the coefficients:
@@
We can extract the model parameters in (10) by fitting it to the coefficients shown in Fig. 2. The results for the parameters, found using a least squares fitting procedure, are shown in Fig. 3. We find $\gamma$ crosses zero around $W\approx 16$, consistent with the accepted $W_{c}\approx 16.5$ and we otherwise have $\gamma\leq 0$ in the extended regime. In addition, we find $\gamma\approx\frac{1}{2d}$ deep in the localized phase, which we predict below in Section III.3.